# Extending Graph Burning to Hypergraphs

### v1

@@ 1 Introduction
We therefore formulate the following rule for how the fire propagates. Fire spreads to a vertex $v$ in round $r$ if and only if there is a non-singleton edge $\{v,u_{1},\ldots,u_{k}\}$ such that each of $u_{1},u_{2},\ldots,u_{k}$ was on fire at the end of round $r-1$. Clearly this reduces to the original game when the hypergraph has only edges of size two. Rounds are indexed by $\mathbb{N}$ starting at one, they have the same structure as in graph burning, and the definitions of a source, redundant source, burning sequence, and burning number are all analogous to those in graph burning.
@@ 2 General Results and Bounds
The bounds in Corollary 2.2, Theorem 2.3, and Corollary 2.5 are tight. Consider the simple hypergraph $H$ in Figure 1. A minimum lazy burning set is $\{x,y,z\}$, an optimal burning sequence is $(x,z,y)$, and $|V(H)|-|E(H)|=3$. Hence, $|V(H)|-|E(H)|=b_{L}(H)=b(H)$, so the aforementioned bounds are all tight simultaneously. This example generalizes to an infinite family of hypergraphs that show tightness; one may construct such a hypergraph $H$ with $V(H)=\{v_{1},\ldots,v_{n}\}$ and $E(H)=\big{\{}\{v_{1},\ldots,v_{n-1}\}\big{\}}$, where $n\geq 3$. Then a minimum lazy burning set is $\{v_{2},\ldots,v_{n}\}$, an optimal burning sequence is $(v_{2},\ldots,v_{n})$, and $|V(H)|-|E(H)|=n-1$.
$x$$y$$z$$w$
Figure 1: An example where all the bounds in Corollary 2.2, Theorem 2.3, and Corollary 2.5 are tight simultaneously.
@@
If $H$ has no isolated vertices then $b_{L}(H)<b(H)$.
@@
Each inequality in Theorem 2.14 is tight, and Figure 2 shows an example of a hypergraph $H$ where all of them are tight simultaneously. It has $7=|V(H)|-|E(H)|=b_{L}(H)$ and $18=b(H)=\alpha(H)+$ 1. Of course, this example can be expanded to an infinite family of hypergraphs that exhibit tightness for each inequality in Theorem 2.14 simultaneously. Simply consider the family of hypergraphs that consist of a single edge containing all of their vertices (excluding the hypergraph which is a single vertex in an edge).
@@
If $V(H)=\{v_{1},\ldots,v_{n}\}$ and $E(H)=\big{\{}\{v_{1},\dots,v_{k}\},\{v_{2},\ldots,v_{k+1}\},\ldots,\{v_{n-k+1},\ldots,v_{n}\}\big{\}}$ then $H$ is called a $k$-uniform tight path. This family of hypergraphs is used to prove Theorem 2.18, which states that the difference between $b_{L}(H)$ and $b(H)$ can be arbitrarily large.
@@
Given a $k$-uniform tight path $H$ on $n$ vertices, define a seed as a set of $k-1$ vertices all belonging to a common edge. Define a burned seed as a seed whose $k-1$ vertices were chosen as sources in the arsonist’s burning sequence.
@@
Given a hypergraph $G$, define $S_{G}(r)$ as the maximum number of vertices that could possibly be on fire in $G$ at the end of round $r$.
@@
Write $V(H)=\{u_{1},\ldots,u_{n}\}$. Clearly, no fire will propagate in $H$ without the existence of a burned seed. Also, once a burned seed $\{u_{i},u_{i+1}\}$ is created in round $r$, it will cause fire to spread to $u_{i-1}$ and $u_{i+2}$ in round $r+1$ (if these vertices are not already on fire). In round $r+2$, it will cause fire to spread to $u_{i-2}$ and $u_{i+3}$, etc. The arsonist’s best strategy is to create a burned seed every second round, spaced such that no vertex $u_{j}$ will catch fire due to propagation from two burned seeds until possibly the final round when $H$ is completely burned. By following this strategy, the maximum number of vertices that can catch fire through propagation in round $r$ is twice the number of burned seeds that existed at the end of round $r-1$. Define $n(r)$ as the number of vertices that become burned in round $r$ when following this strategy (including sources), so $n(r)=S_{H}(r)-S_{H}(r-1)$. Clearly $n(1)=n(2)=1$ since no propagation can occur in the first two rounds. There is one burned seed at the start of rounds three and four, so $n(3)=n(4)=2(1)+1=3$. Similarly, there are two burned seeds at the start of rounds five and six, so $n(5)=n(6)=2(2)+1=5$. The sequence of $n(r)$-values is therefore $1,1,3,3,5,5,7,7,\ldots$ Observe that $S_{H}(r)=\sum_{i=1}^{r}n(i)$, so the sequence of $S_{H}(r)$-values is $1,2,5,8,13,18,25,32,41,50,\ldots$ The formula for this sequence is known to be $S_{H}(r)=\left\lfloor\frac{r^{2}+1}{2}\right\rfloor$; see sequence A000982 in [21].
@@ 3 Disconnected Hypergraphs
In this section we consider how to write the (lazy) burning number of a disconnected hypergraph in terms of the (lazy) burning numbers of its connected components. The solution is trivial in the lazy case (see Lemma 3.1), but the round-based case is more complicated.
@@
The upper bound from Lemma 3.3 is tight. Consider the hypergraph $H$ with $V(H)=\{1,2,.\ldots,21\}$ and $E(H)=\big{\{}\{1,2,3\},\{4,5,6\},\ldots,\{19,20,21\}\big{\}}$. The subhypergraph induced by one of the seven edges has burning number $3$, and since (1,2,4,5,7,8,10,11,13,14,(1,2,4,5,7,8,10,11,13,14,( 1 , 2 , 4 , 5 , 7 , 8 , 10 , 11 , 13 , 14 , 16,17,19,20,21)16,17,19,20,21)16 , 17 , 19 , 20 , 21 ) is an optimal burning sequence for $H$, it has burning number $15=(3)(7)-7+1$. Indeed, this example can be expanded to an infinite family of hypergraphs which exhibit the tightness of the bound in Lemma 3.3. Such a hypergraph $G$ consists of multiple disjoint non-singleton edges $e_{1},\ldots,e_{k}$, with no other edges and no isolated vertices. For each $i\in\{1,\ldots,k\}$, let $G_{i}$ be the subhypergraph of $G$ induced by $e_{i}$. Then, one may construct an optimal burning sequence for $G$ by burning $|e_{i}|-1=b(G_{i})-1$ sources in $e_{i}$ for each $i\in\{1,\ldots,k-1\}$, followed by $|e_{k}|=b(G_{k})$ sources in $e_{k}$ (where the last source is redundant). This shows the tightness of the bound in Lemma 3.3, since the total number of sources is
@@
There are also examples where the upper bound from Lemma 3.3 is a strict inequality. Consider the hypergraph in Figure 6. The burning numbers of the left, middle, and right connected components are $5$, $3$, and $2$ respectively. But $(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6})$ is an optimal burning sequence for the entire hypergraph, so it has burning number $6<(5+3+2)-3+1$. It is an open question as to whether or not there is an infinite family of hypegraphs for which the bound in Lemma 3.3 is strict.
@@ 4 Subhypergraphs
In this section we compare the (lazy) burning number of a hypergraph $H$ to the (lazy) burning numbers of various types of subhypergraphs of $H$. We now briefly define the subhypergraphs of interest.
A hypergraph $H^{\prime}=(V^{\prime},E^{\prime})$ is called a weak subhypergraph (or just subhypergraph) of $H=(V,E)$ if $V^{\prime}\subseteq V$ and either $E^{\prime}=\emptyset$, or, after a suitable permutation of its rows and columns, the incidence matrix for $H^{\prime}$ is a submatrix of the incidence matrix of $H$. Thus, each edge $e^{\prime}\in E^{\prime}$ has $e^{\prime}=e\cap V^{\prime}$ for some $e\in E$ (so $H^{\prime}$ may not be simple). If, in addition, $E^{\prime}=\{e\cap V^{\prime}\mid e\in E,e\cap V^{\prime}\neq\emptyset\}$, then $H^{\prime}$ is said to be induced by $V^{\prime}$ (that is, all nonempty edges $e\cap V^{\prime}$, singletons included, are present in $E^{\prime}$).
A hypergraph $H^{\prime\prime}=(V^{\prime\prime},E^{\prime\prime})$ is called a strong subhypergraph (or hypersubgraph) of $H=(V,E)$ if $V^{\prime\prime}\subseteq V$ and $E^{\prime\prime}\subseteq E$. If all edges induced by $V^{\prime\prime}$ (i.e. all $e\subseteq V^{\prime\prime}$) are present in $E^{\prime\prime}$ then $H^{\prime\prime}$ is said to be induced by $V^{\prime\prime}$, and we write $H^{\prime\prime}=H[V^{\prime\prime}]$. If $V^{\prime\prime}=\cup_{e\in E^{\prime\prime}}e$ then $H^{\prime\prime}$ is said to be induced by $E^{\prime\prime}$, and we write $H^{\prime\prime}=H[E^{\prime\prime}]$. Note that every strong subhypergraph of $H$ is also a weak subhypergraph of $H$, but the converse is not necessarily true.
@@
We must introduce some more notation for use in the following proof. In the round-based version of the game, denote the set of vertices that are on fire in a hypergraph $H$ at the end of round $r$ when burning according to the sequence $S$ by $F(H,S,r)$.
@@
    Otherwise, if $u_{r}\in V(G)$ but $u_{r}$ was on fire at the end of round $r-1$, then choose any unburned vertex in $V(G)$ as the $r^{th}$ source in $S^{\prime}$.
@@
    Otherwise, if $u_{r}\in V(G)$ and $u_{r}$ was not on fire at the end of round $r-1$, then choose $u_{r}$ as the $r^{th}$ source in $S^{\prime}$.
Clearly, by the above construction, $|S^{\prime}|\leq|S|$. Moreover, $|S^{\prime}|<|S|$ is only true if (1) occurs, in which case $S^{\prime}$ is indeed a burning sequence for $G$. We therefore assume that (1) does not occur, and hence $|S^{\prime}|=|S|$. We must show that $S^{\prime}$ leaves $G$ fully burned after the final round, $b(H)$. In particular, we show $F(H,S,r)\cap V(G)\subseteq F(G,S^{\prime},r)$ for each round $r\in\{1,2,\ldots,b(H)\}$ by induction.
Base case. Consider the earliest round $r_{0}$ at which $F(H,S,r_{0})\cap V(G)\neq\emptyset$, since the inclusion clearly holds for all earlier rounds. Recall that every edge in $H$ contains at least two vertices of $G$. Hence, for any fire to spread to a vertex of $G$ in $BG(H,S)$, a vertex of $G$ must first be chosen as a source in $S$ at a strictly earlier round. Hence, $F(H,S,r_{0})\cap V(G)=\{u_{r_{0}}\}$ (i.e. the first vertex of $G$ to be on fire in $BG(H,S)$ is a source). But by the above construction, in $BG(G,S^{\prime})$, either $u_{r_{0}}$ is chosen as the $r_{0}^{th}$ source, or it was on fire in a strictly earlier round. Therefore, $F(H,S,r_{0})\cap V(G)=\{u_{r_{0}}\}\subseteq F(G,S^{\prime},r_{0})$.
Inductive hypothesis. Suppose $F(H,S,r)\cap V(G)\subseteq F(G,S^{\prime},r)$ for some round $r\geq r_{0}$.
Inductive step. By the construction of $S^{\prime}$, if $u_{r+1}\in V(G)$ then $u_{r+1}$ is on fire in BG$(G,S^{\prime})$ at the end of round $r+1$. Consider any non-source vertex $v\in V(G)$ that catches fire through propagation in $BG(H,S)$ in round $r+1$. If $v$ was on fire in $BG(G,S^{\prime})$ at an earlier round then there is nothing to show, so assume the contrary. We must show that $v$ catches fire through propagation in $BG(G,S^{\prime})$ in round $r+1$.
Consider an edge $e\in E(H)$ that caused fire to spread to $v$ in $BG(H,S)$. At the end of round $r$, each vertex in $e\setminus\{v\}$ was on fire. Recall that $e\cap V(G)$ is a non-singleton edge in $G$. By the inductive hypothesis, $F(H,S,r)\cap V(G)\subseteq F(G,S^{\prime},r)$, and in particular, $(e\cap V(G))\setminus\{v\}\subseteq F(H,S,r)\cap V(G)\subseteq F(G,S^{\prime},r)$. Therefore, each vertex of $(e\cap V(G))\setminus\{v\}$ was on fire in $BG(G,S^{\prime})$ at the end of round $r$. But then the edge $e\cap V(G)$ causes fire to spread to $v$ in $BG(G,S^{\prime})$ in round $r+1$.
Hence, $F(H,S,r+1)\cap V(G)\subseteq F(G,S^{\prime},r+1)$. We therefore have $V(G)=F(H,S,b(H))\cap V(G)\subseteq F(G,S^{\prime},b(H))$, so $F(G,S^{\prime},b(H))=V(G)$. That is, $S^{\prime}$ is a burning sequence for $G$, so $b(G)\leq|S^{\prime}|=|S|=b(H)$. ∎
@@
Now, let $G_{1}=(V_{1},E_{1})$ where $V_{1}=\{u_{3},u_{4},\ldots,u_{n}\}$ and $E_{1}=\big{\{}\{u_{3},u_{4},\ldots,u_{n}\}\big{\}}$, so $G_{1}$ is strongly induced by $u_{3},u_{4},\ldots,u_{n}$. Clearly, $b(G_{1})=n-2$ and $b_{L}(G_{1})=n-3$.
Finally, let $G_{2}=(V_{2},E_{2})$ where $V_{2}=\{u_{3},u_{4},\ldots,u_{n-1}\}$ and $E_{2}=\big{\{}\{u_{3}\},\{u_{4}\},\ldots,\{u_{n-1}\},$$\{u_{3},u_{4},\ldots,u_{n-1}\}\big{\}}$, so $G_{2}$ is weakly induced by $u_{3},u_{4},\ldots,u_{n-1}$. Clearly, $b(G_{2})=n-3$ and $b_{L}(G_{2})=n-4$. ∎
@@
Notice that the weak induced subhypergraph $G_{2}$ of $H$ in Lemma 4.5 need not contain fewer edges than $H$. That is, if instead $G_{2}$ was induced by $u_{3},\ldots,u_{n}$, it would still have a larger (lazy) burning number than $H$. It does however contain singleton edges, and we are unsure if this must always be the case. That is, we would like to determine if there exists a class of hypergraphs $H$ with weak induced subhypergraphs $G_{2}$ such that $b(G_{2})-b(H)$ and $b_{L}(G_{2})-b_{L}(H)$ are both unbounded, and $G_{2}$ contains no singleton edges. Of course, in light of Lemmas 4.3 and 4.4, such a hypergraph $G_{2}$ would need to have fewer edges than $H$.

### v2

@@ 1 Introduction
We therefore formulate the following rule for how the fire propagates. Fire spreads to a vertex $v$ in round $r$ if and only if there is a non-singleton edge $\{v,u_{1},\ldots,u_{k}\}$ such that each of $u_{1},u_{2},\ldots,u_{k}$ was on fire at the end of round r1r-1. Clearly this reduces to the original game when the hypergraph has only edges of size two. Rounds are indexed by $\mathbb{N}$ starting at one, they have the same structure as in graph burning, and the definitions of a source, redundant source, burning sequence, and burning number are all analogous to those in graph burning.
@@ 2 General Results and Bounds
The bounds in Corollary 2.2, Theorem 2.3, and Corollary 2.5 are tight. Consider the simple hypergraph $H$ in Figure 1. A minimum lazy burning set is $\{x,y,z\}$, an optimal burning sequence is $(x,z,y)$, and |V(H)||E(H)|=3|V(H)|-|E(H)|=3. Hence, $|V(H)|-|E(H)|=b_{L}(H)=b(H)$, so the aforementioned bounds are all tight simultaneously. This example generalizes to an infinite family of hypergraphs that show tightness; one may construct such a hypergraph $H$ with $V(H)=\{v_{1},\ldots,v_{n}\}$ and $E(H)=\big{\{}\{v_{1},\ldots,v_{n-1}\}\big{\}}$, where $n\geq 3$. Then a minimum lazy burning set is $\{v_{2},\ldots,v_{n}\}$, an optimal burning sequence is $(v_{2},\ldots,v_{n})$, and |V(H)||E(H)|=n1|V(H)|-|E(H)|=n-1.
xxyyzzww
Figure 1: An example of a hypergraph $H$ where all the bounds in Corollary 2.2, Theorem 2.3, and Corollary 2.5 are tight simultaneously. Note that $|E(H)|=1$ since singleton edges are not allowed in simple hypergraphs.
@@
If $|V(H)|\geq 2$ and $H$ has no isolated vertices then $b_{L}(H)<b(H)$.
@@
The difference between $b_{L}(H)$ and $\alpha(H)$ (respectively $b(H)$ and $\alpha(H)+1$) can also be arbitrarily large. Consider the family of hypergraphs $H$ with $V(H)=\{v_{1},\ldots,v_{n}\}$ and $E(H)=\big{\{}\{v_{1},v_{2},v_{3}\},\{v_{1},v_{2},v_{4}\},\ldots,\{v_{1},v_{2},v_{n}\}\big{\}}$. Such a hypergraph has $b_{L}(H)=2$, $b(H)=3$, and $\alpha(H)=n-1$.
@@
Each inequality in Theorem 2.14 is tight, and Figure 2 shows an example of a hypergraph $H$ where all of them are tight simultaneously. It has $7=|V(H)|-|E(H)|=b_{L}(H)$ and $8=b(H)=\alpha(H)+1$. Of course, this example can be expanded to an infinite family of hypergraphs that exhibit tightness for each inequality in Theorem 2.14 simultaneously. Simply consider the family of hypergraphs that consist of a single edge containing all of their vertices (excluding the hypergraph which is a single vertex in an edge).
@@
Considering the inequalities in Theorem 2.14 further, we have already shown that the difference between $\alpha(H)+1$ and $b(H)$ can be arbitrarily large. Additionally it is easy to see that $b_{L}(H)-(|V(H)|-|E(H)|)$ can also be arbitrarily large, by noting that $b_{L}(H)>0$ for every hypergraph $H$, but, for instance, taking $H$ to be the $k$-uniform complete hypergraph of order $n$, we have that $|V(H)|-|E(H)|=n-\binom{n}{k}$.
To see that the difference between the burning number $b(H)$ and the lazy burning number $b_{L}(H)$ can be arbitrarily large (Corollary 2.18), we will consider tight paths. Specifically, the $k$-uniform tight path of order $n$ is the hypergraph $H$ with $V(H)=\{v_{1},\ldots,v_{n}\}$ and
$E(H)=\big{\{}\{v_{1},\dots,v_{k}\},\{v_{2},\ldots,v_{k+1}\},\ldots,\{v_{n-k+1},\ldots,v_{n}\}\big{\}}.$
@@
Given a $k$-uniform tight path $H$ on $n$ vertices, define a seed as a set of k1k-1 vertices all belonging to a common edge. Define a burned seed as a seed whose k1k-1 vertices were chosen as sources in the arsonist’s burning sequence.
@@
Given a hypergraph $G$, define $S_{G}(r)$ as the maximum number of vertices that could possibly be on fire in $G$ at the end of round $r$ over all possible burning sequences. Note that the definition of $S_{G}(r)$ only applies to the round-based model of hypergraph burning.
@@
Write $V(H)=\{u_{1},\ldots,u_{n}\}$. Clearly, no fire will propagate in $H$ without the existence of a burned seed. Also, once a burned seed $\{u_{i},u_{i+1}\}$ is created in round $r$, it will cause fire to spread to $u_{i-1}$ and $u_{i+2}$ in round $r+1$ (if these vertices are not already on fire). In round $r+2$, it will cause fire to spread to $u_{i-2}$ and $u_{i+3}$, etc. The arsonist’s best strategy is to create a burned seed every second round, spaced such that no vertex $u_{j}$ will catch fire due to propagation from two burned seeds until possibly the final round when $H$ is completely burned. By following this strategy, the maximum number of vertices that can catch fire through propagation in round $r$ is twice the number of burned seeds that existed at the end of round r1r-1. Define $n(r)$ as the number of vertices that become burned in round $r$ when following this strategy (including sources), so $n(r)=S_{H}(r)-S_{H}(r-1)$. Clearly $n(1)=n(2)=1$ since no propagation can occur in the first two rounds. There is one burned seed at the start of rounds three and four, so $n(3)=n(4)=2(1)+1=3$. Similarly, there are two burned seeds at the start of rounds five and six, so $n(5)=n(6)=2(2)+1=5$. The sequence of $n(r)$-values is therefore $1,1,3,3,5,5,7,7,\ldots$ . Observe that $S_{H}(r)=\sum_{i=1}^{r}n(i)$, so the sequence of $S_{H}(r)$-values is $1,2,5,8,13,18,25,32,41,50,\ldots$ . The formula for this sequence is known to be $S_{H}(r)=\left\lfloor\frac{r^{2}+1}{2}\right\rfloor$; see sequence A000982 in [23].
@@
We will describe the optimal burning sequence for a tight 3-uniform path $H_{n}$ on $n$ vertices. First, we recall how to find the optimal burning sequence on a path $P_{n}$ from [8], as this is the inspiration for how we burn $H_{n}$. In particular, $b(P_{n})=\left\lceil\sqrt{n}\ \right\rceil$ [8], and if $n$ is a square number, $P_{n}$ can be burned “perfectly” (i.e. with no overlap of propagating fires) in $\sqrt{n}$ rounds as follows. Write $V(P_{n})=\{u_{1},\ldots,u_{n}\}$ and $E(P_{n})=\big{\{}\{u_{1},u_{2}\},\ldots,\{u_{n-1},u_{n}\}\big{\}}$. Set $k=\sqrt{n}$, and denote the burning sequence by $(x_{1},\ldots,x_{k})$. Finally, for $i=0,1,\ldots,k-1$, set $x_{k-i}=u_{n-i^{2}-i}$. If $n$ is not a square number, let $N$ be the smallest square number exceeding $n$, so $b(P_{n})=b(P_{N})$. Consider an optimal burning sequence on $P_{N}$ as described above. Then, delete the NnN-n lowest-indexed vertices $u_{i}$ to create $P_{n}$ from $P_{N}$. If the first source $x_{1}$ was deleted, the new first source becomes $u_{1+N-n}$ (i.e. the new lowest-indexed vertex in the resulting path).
Our strategy for burning $H_{n}$ is similar. If $n$ is twice a square number or $\sqrt{2n-1}\in\mathbb{N}$, then $H_{n}$ can be burned “perfectly” in $\left\lceil\sqrt{2n-1}\ \right\rceil$ rounds as follows. Write $V(H_{n})=\{u_{1},\ldots,u_{n}\}$ and $E(H_{n})=\big{\{}\{u_{1},u_{2},u_{3}\},\ldots,\{u_{n-2},u_{n-1},u_{n}\}\big{\}}$. Set $k=\left\lceil\sqrt{2n-1}\ \right\rceil$ and denote the burning sequence by $(x_{1},\ldots,x_{k})$. Note that if $n$ is twice a square number then $k$ will be even, and if $\sqrt{2n-1}\in\mathbb{N}$ then $k$ will be odd. Now, for $s=1,2,\ldots,\left\lceil\frac{k}{2}\right\rceil$, the $s^{th}$ odd-indexed source is $x_{2s-1}=u_{a}$, where
$a=k-2s+2+2\sum_{i=1}^{s-1}(k-2i+1)\ .$
Similarly, for $s=1,2,\ldots,\left\lfloor\frac{k}{2}\right\rfloor$, the $s^{th}$ even-indexed source is $x_{2s}=u_{b}$, where
$b=k-2s+1+2\sum_{i=1}^{s-1}(k-2i+1)\ .$
Of course, $x_{2s-1}$ and $x_{2s}$ make up the $s^{th}$ seed. These formulas were found using a technique similar to the one used in [8]. In particular, the $s^{th}$ seed causes all vertices within distance k2sk-2s to catch fire by the end of round $k$. Spacing the seeds “perfectly” so that there is no overlap in propagating fires gives the above formulas, as well as the fact that $n$ is either twice a square number or $\sqrt{2n-1}\in\mathbb{N}$. Finally, if $n$ has neither of these properties, then let $N$ be the smallest number exceeding $n$ that is twice a square number or satisfies $\sqrt{2N-1}\in\mathbb{N}$. Then $b(H_{n})=b(H_{N})$. Consider an optimal burning sequence on $H_{N}$ as described above, then delete the NnN-n lowest-indexed vertices $u_{i}$ to create $H_{n}$ from $H_{N}$. If $x_{1}$ and/or $x_{2}$ were deleted, then the new first seed becomes $u_{1+N-n}$ and $u_{2+N-n}$.
@@ 3 Disconnected Hypergraphs
In this section we consider how to write the burning number of a disconnected hypergraph in terms of the burning numbers of its connected components. The solution is trivial in the lazy case (see Lemma 3.1), but the round-based case is more complicated.
@@
The upper bound from Lemma 3.3 is tight. Consider the hypergraph $H$ with $V(H)=\{1,2,.\ldots,21\}$ and $E(H)=\big{\{}\{1,2,3\},\{4,5,6\},\ldots,\{19,20,21\}\big{\}}$. The subhypergraph induced by one of the seven edges has burning number $3$, and since $(1,2,4,5,7,8,10,11,13,14,$ $16,17,19,20,21)$ is an optimal burning sequence for $H$, it has burning number 15=(3)(7)7+115=(3)(7)-7+1. Indeed, this example can be expanded to an infinite family of hypergraphs which exhibit the tightness of the bound in Lemma 3.3. Such a hypergraph $G$ consists of multiple disjoint non-singleton edges $e_{1},\ldots,e_{k}$, with no other edges and no isolated vertices. For each $i\in\{1,\ldots,k\}$, let $G_{i}$ be the subhypergraph of $G$ induced by $e_{i}$. Then, one may construct an optimal burning sequence for $G$ by burning $|e_{i}|-1=b(G_{i})-1$ sources in $e_{i}$ for each $i\in\{1,\ldots,k-1\}$, followed by $|e_{k}|=b(G_{k})$ sources in $e_{k}$ (where the last source is redundant). This shows the tightness of the bound in Lemma 3.3, since the total number of sources is
@@
There are also examples where the upper bound from Lemma 3.3 is a strict inequality. Consider the hypergraph in Figure 6. The burning numbers of the left, middle, and right connected components are $5$, 33, and $2$ respectively. But $(u_{1},u_{2},u_{3},u_{4},u_{5},u_{6})$ is an optimal burning sequence for the entire hypergraph, so it has burning number 6<(5+3+2)3+16<(5+3+2)-3+1. It is an open question as to whether or not there is an infinite family of hypegraphs for which the bound in Lemma 3.3 is strict.
@@ 4 Subhypergraphs
In this section we compare the (lazy) burning number of a hypergraph $H$ to the (lazy) burning numbers of various types of subhypergraphs of $H$. Before describing the subhypergraphs of interest, we first define the incidence matrix of a hypergraph $H=(V,E)$, which is the $|V|\times|E|$ matrix whose $(i,j)$-entry is a 1 if the $i^{th}$ vertex belongs to the $j^{th}$ edge, and a 0 otherwise.
A hypergraph $H^{\prime}=(V^{\prime},E^{\prime})$ is called a weak subhypergraph (or just subhypergraph) of $H=(V,E)$ if $V^{\prime}\subseteq V$ and either $E^{\prime}=\emptyset$, or, after a suitable permutation of its rows and columns, the incidence matrix for $H^{\prime}$ is a submatrix of the incidence matrix of $H$. Thus, each edge $e^{\prime}\in E^{\prime}$ has $e^{\prime}=e\cap V^{\prime}$ for some $e\in E$ (so $H^{\prime}$ may not be simple, even if $H$ is). If, in addition, $E^{\prime}=\{e\cap V^{\prime}\mid e\in E,e\cap V^{\prime}\neq\emptyset\}$, then $H^{\prime}$ is said to be induced by $V^{\prime}$ (that is, all nonempty edges $e\cap V^{\prime}$, singletons included, are present in $E^{\prime}$).
A hypergraph $H^{\prime\prime}=(V^{\prime\prime},E^{\prime\prime})$ is called a strong subhypergraph (or hypersubgraph) of $H=(V,E)$ if $V^{\prime\prime}\subseteq V$ and $E^{\prime\prime}\subseteq E$. If all edges $e\in E$ such that $e\subseteq V^{\prime\prime}$ are present in $E^{\prime\prime}$, then $H^{\prime\prime}$ is said to be induced by $V^{\prime\prime}$, and we write $H^{\prime\prime}=H[V^{\prime\prime}]$. If $V^{\prime\prime}=\cup_{e\in E^{\prime\prime}}e$ then $H^{\prime\prime}$ is said to be induced by $E^{\prime\prime}$, and we write $H^{\prime\prime}=H[E^{\prime\prime}]$. Note that every strong subhypergraph of $H$ is also a weak subhypergraph of $H$, but the converse is not necessarily true.
@@
We must introduce some more notation for use in the following proof. In the round-based version of the game, denote the set of vertices that are on fire in a hypergraph $H$ at the end of round $r$ when burning according to the sequence $S$ by $F_{r}(H,S)$.
@@
    Otherwise, if $u_{r}\in V(G)$ but $u_{r}$ was on fire at the end of round r1r-1, then choose any unburned vertex in $V(G)$ as the $r^{th}$ source in $S^{\prime}$.
@@
    Otherwise, if $u_{r}\in V(G)$ and $u_{r}$ was not on fire at the end of round r1r-1, then choose $u_{r}$ as the $r^{th}$ source in $S^{\prime}$.
Clearly, by the above construction, $|S^{\prime}|\leq|S|$. Moreover, $|S^{\prime}|<|S|$ is only true if (1) occurs, in which case $S^{\prime}$ is indeed a burning sequence for $G$. We therefore assume that (1) does not occur, and hence $|S^{\prime}|=|S|$. We must show that $S^{\prime}$ leaves $G$ fully burned after the final round, $b(H)$. In particular, we show $F_{r}(H,S)\cap V(G)\subseteq F_{r}(G,S^{\prime})$ for each round $r\in\{1,2,\ldots,b(H)\}$ by induction.
Base case. Consider the earliest round $r_{0}$ at which $F_{r_{0}}(H,S)\cap V(G)\neq\emptyset$, since the inclusion clearly holds for all earlier rounds. Recall that every edge in $H$ contains at least two vertices of $G$. Hence, for any fire to spread to a vertex of $G$ in $BG(H,S)$, a vertex of $G$ must first be chosen as a source in $S$ at a strictly earlier round. Hence, $F_{r_{0}}(H,S)\cap V(G)=\{u_{r_{0}}\}$ (i.e. the first vertex of $G$ to be on fire in $BG(H,S)$ is a source). But by the above construction, in $BG(G,S^{\prime})$, either $u_{r_{0}}$ is chosen as the $r_{0}^{th}$ source, or it was on fire in a strictly earlier round. Therefore, $F_{r_{0}}(H,S)\cap V(G)=\{u_{r_{0}}\}\subseteq F_{r_{0}}(G,S^{\prime})$.
Inductive hypothesis. Suppose $F_{r}(H,S)\cap V(G)\subseteq F_{r}(G,S^{\prime})$ for some round $r\geq r_{0}$.
[…]
Consider an edge $e\in E(H)$ that caused fire to spread to $v$ in $BG(H,S)$. At the end of round $r$, each vertex in $e\setminus\{v\}$ was on fire. Recall that $e\cap V(G)$ is a non-singleton edge in $G$. By the inductive hypothesis, $F_{r}(H,S)\cap V(G)\subseteq F_{r}(G,S^{\prime})$, and in particular, $(e\cap V(G))\setminus\{v\}\subseteq F_{r}(H,S)\cap V(G)\subseteq F_{r}(G,S^{\prime})$. Therefore, each vertex of $(e\cap V(G))\setminus\{v\}$ was on fire in $BG(G,S^{\prime})$ at the end of round $r$. But then the edge $e\cap V(G)$ causes fire to spread to $v$ in $BG(G,S^{\prime})$ in round $r+1$.
Hence, $F_{r+1}(H,S)\cap V(G)\subseteq F_{r+1}(G,S^{\prime})$. We therefore have $V(G)=F_{b(H)}(H,S)\cap V(G)\subseteq F_{b(H)}(G,S^{\prime})$, so $F_{b(H)}(G,S^{\prime})=V(G)$. That is, $S^{\prime}$ is a burning sequence for $G$, so $b(G)\leq|S^{\prime}|=|S|=b(H)$. ∎
@@
Now, let $G_{1}=(V_{1},E_{1})$ where $V_{1}=\{u_{3},u_{4},\ldots,u_{n}\}$ and $E_{1}=\big{\{}\{u_{3},u_{4},\ldots,u_{n}\}\big{\}}$, so $G_{1}$ is strongly induced by $\{u_{3},u_{4},\ldots,u_{n}\}$. Clearly, $b(G_{1})=n-2$ and $b_{L}(G_{1})=n-3$.
Finally, let $G_{2}=(V_{2},E_{2})$ where $V_{2}=\{u_{3},u_{4},\ldots,u_{n-1}\}$ and $E_{2}=\big{\{}\{u_{3}\},\{u_{4}\},\ldots,\{u_{n-1}\},$ $\{u_{3},u_{4},\ldots,u_{n-1}\}\big{\}}$, so $G_{2}$ is weakly induced by $\{u_{3},u_{4},\ldots,u_{n-1}\}$. Clearly, $b(G_{2})=n-3$ and $b_{L}(G_{2})=n-4$. ∎
@@
Notice that the weak induced subhypergraph $G_{2}$ of $H$ in Lemma 4.5 need not contain fewer edges than $H$. That is, if instead $G_{2}$ was induced by $\{u_{3},\ldots,u_{n}\}$, it would still have a larger (lazy) burning number than $H$. It does however contain singleton edges, and we are unsure if this must always be the case. That is, we would like to determine if there exists a class of hypergraphs $H$ with weak induced subhypergraphs $G_{2}$ such that $b(G_{2})-b(H)$ and $b_{L}(G_{2})-b_{L}(H)$ are both unbounded, and $G_{2}$ contains no singleton edges. Of course, in light of Lemmas 4.3 and 4.4, such a hypergraph $G_{2}$ would need to have fewer edges than $H$.